# TIMeSynC: Temporal Intent Modelling with Synchronized Context Encodings for Financial Service Applications

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2024conference: Eighteenth ACM Conference on Recommender Systems; Oct 14–18, 2024; Bari, Italyccs: Computing methodologies Machine learningccs: Information systems Recommender systemsccs: Computing methodologies Knowledge representation and reasoningccs: Information systems Temporal data
@@ 2.1. Sequential Recommendation
2.2. Handling Tabular Data with LLM
@@ 3. Problem Statement
The sequential intent prediction problem in finance aims to forecast user actions based on inputs from diverse domains. Let $X=[X_{1},X_{2},\ldots,X_{n}]$ represent $n$ domains, where $X_{i}$ denotes different types of financial interactions, such as credit card/debit card transactions, payments, rewards, account status, customer care messages, call center intents, digital page views, etc. Each $X_{i}$ is a time series represented as $[X_{i1},X_{i2},\ldots,X_{it}]$. The goal is to predict future user’s intents $Y=[y_{1},y_{2},\ldots,y_{k}]$ at time $t$, given the financial context $X$ and the prior sequence of user’s intents $Y$.
@@ 4.1. Experimental Setup
[…]
@@ 4.2.2. TIMeSynC:
[…]

### v4

@@ Abstract.
copyright: acmlicensedjournalyear: 2024conference: Eighteenth ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italyccs: Computing methodologies Machine learningccs: Information systems Recommender systemsccs: Computing methodologies Knowledge representation and reasoningccs: Information systems Temporal data
@@ 2.1. Sequential Recommendation
$2.$ Handling Tabular Data with LLM
@@ 3. Problem Statement
The sequential intent prediction problem in finance aims to forecast user actions based on inputs from diverse domains. Let $X=[X_{1},X_{2},\ldots,X_{n}]$ represent $n$ domains, where $X_{i}$ denotes different types of financial interactions, such as credit card/debit card transactions, payments, rewards, account status, customer care messages, call center intents, digital page views, etc. Each $X_{i}$ is a time series represented as $X_{i1},X_{i2},\ldots,X_{it}]$. The goal is to predict future user’s intents $Y=[y_{1},y_{2},\ldots,y_{k}]$ at time $t$, given the financial context $X$ and the prior sequence of user’s intents $Y$.
@@ 4.1. Experimental Setup
$\text{Recall@k}(u)=\frac{\left|\text{Intents}(u)\cap\text{Top}_{k}(\text{Ranked\_Intents}(u))\right|}{\left|\text{Intents}(u)\right|}$
@@ 4.2.2. TIMeSynC:
  • Intent Decoder: The Transformer Decoder Layer, as described in (Vaswani, 2017), processes a raw sequence of intents. We modify the decoder layer as follows: The cross-attention mechanism uses Queries ($q_{i}$) from the intent self-attention, Keys ($k_{j}$), and Values ($v_{j}$) from the output of the context encoder ($E$). TimeAliBi ($D(t)$, $E(t)$) is applied to represent the relativity of time, followed by a causal mask $\text{causal\_mask}(D(t),E(t))$, where $D(t)$ denotes the time associated with the decoder sequence. Similar to the encoder, time representations are also incorporated in the decoder as shown in Figure 1.